A fountain propels 30 000kg of water per minute to a height of 140m. Calculate the power used in raising the water.
1
Expert's answer
2018-01-29T10:14:08-0500
dm/dt=30000 kg/min=30000/60 kg/s=500 kg/s
The power used in raising the water is
P=dW/dt=d/dt (mgh)=dm/dt gh.
P=(500)(10)(140)=700000 W=700 kW.